# Neural Context Flows for Meta-Learning of Dynamical Systems

### v4

@@ Abstract
Neural Ordinary Differential Equations (NODEs) often struggle to adapt to new dynamic behaviours caused by parameter changes in the underlying system, even when these dynamics are similar to previously observed behaviours. This problem becomes more challenging when the changing parameters are unobserved, meaning their value or influence cannot be directly measured when collecting data. To address this issue, we introduce Neural Context Flow (NCF), a robust and interpretable Meta-Learning framework that includes uncertainty estimation. NCF uses higher-order Taylor expansion to enable contextual self-modulation, allowing context vectors to influence dynamics from other domains while also modulating themselves. After establishing convergence guarantees, we empirically test NCF and compare it to related adaptation methods. Our results show that NCF achieves state-of-the-art Out-of-Distribution performance on 5 out of 6 linear and non-linear benchmark problems. Through extensive experiments, we explore the flexible model architecture of NCF and the encoded representations within the learned context vectors. Our findings highlight the potential implications of NCF for foundational models in the physical sciences, offering a promising approach to improving the adaptability and generalization of NODEs in various scientific applications. Our code is openly available at https://github.com/ddrous/ncflow.
@@ 1 Introduction
A prototypical autonomous dynamical system describes the continuous change, through time $t\in\mathbb{R}^{+}$, of a quantity $x(t)\in\mathbb{R}^{d}$. Its dynamics are heavily dependent on its parameters $c\in\mathbb{R}^{d_{c}}$, according to the (ordinary) differential equation
@@
Consider the problem of reconstructing the mechanical motion of an undamped pendulum, with limited data collected from two domains a.k.a. environments: Earth’s surface, and outer space. Unaware of the physical discrepancies between these environments (gravitational constant, ambient temperature, etc.), we could learn one single environment-agnostic vector field using all available data in a strategy termed One-For-All (OFA) Yin et al. (2021a). However, our model would struggle not only to meaningfully fit such inhomogeneous data, but also to generalize when presented with data from a different environment, e.g., the Moon’s surface. This problem is illustrated in Figure 1 when gravity is the parameter that changes. Similarly, learning one vector field per environment would fail to capture the commonalities between the training environments, which hinders fast adaptation to new ones Yin et al. (2021a). In these conditions, it is imperative to learn what to learn from all the data while acknowledging each environment’s unique properties.
Neural Ordinary Differential Equations (Neural ODEs) Chen et al. (2018) have emerged as a versatile class of methods for learning ordinary, stochastic, and controlled differential equations Kidger et al. (2020). They can be viewed as continuous-time limits of Residual Neural Networks He et al. (2016); Haber and Ruthotto (2017); Weinan (2017), and were popularized in the Dynamical Systems community through Universal Differential Equations Rackauckas et al. (2020). Trained using Differentiable Programming techniques Nzoyem et al. (2023); Ma et al. (2021), they have shown extensive applications in areas like chemical engineering Owoyele and Pal (2022), geosciences Shen et al. (2023), turbulence forecasting Portwood et al. (2019), etc., where the parameters are typically known or the environments fixed.
@@
In Scientific Machine Learning Cuomo et al. (2022) the problem of generalization has largely been tackled by injecting domain knowledge. It is commonly understood that adding a term in Equation 1 that captures as much of the dynamics as possible leads to lower evaluation losses Yin et al. (2021b). For such terms to be added, however, it is essential to have knowledge of the parameters that change, which may then either be directly estimated, or predicted by a neural network within the vector field Rackauckas et al. (2020). We are naturally left to wonder how to efficiently learn a generalizable dynamical system when such physics are absent.
Under constantly changing experimental conditions, two major obstacles to learning the parameter dependence of a vector field arise:
  1. (P1)
    Limited data: models like Neural ODEs are known to be data-hungry Yin et al. (2021b), and limited data from each environment may not be enough to learn a vector field suitable for all environments.
  2. (P2)
    Unobserved parameters: during the data collection process, one might be unfamiliar with the basic physics of the system, or may not know which or how many parameters might be worth paying attention to.
Solving these two problems would contribute to the efficient generalization of the learned models, an urgent concern given the tendency of the Machine Learning community to prioritize Out-of-Distribution (OoD) detection at the expense of adaptation to new settings Arjovsky (2020). OoD adaptation would massively reduce cost and benefit recent efforts towards foundational models in the physical sciences Subramanian et al. (2024); McCabe et al. (2023).
@@
This work presents Neural Context Flows (NCFs), a novel approach for multi-environment generalization of dynamical systems. By leveraging the regularity of the vector field with respect to unobserved parameters, NCFs employ a neural network vector field shared between all environments. This vector field is subsequently modulated by environment-specific latent context vectors. Our contributions are threefold:
@@
    We introduce a robust Meta-Learning methodology for enhancing the generalizability of dynamical systems. This approach effectively addresses problems (P1) and (P2), challenging the prevailing notion that standard Deep Learning techniques are inherently inefficient for Out-of-Distribution (OoD) adaptation Mouli et al. (2024).
@@
    We present an interpretable framework for Multi-Task representation learning, incorporating a straightforward method for uncertainty quantification. This work extends the emerging trend of explainable linearly parameterized physical systems Blanke and Lelarge (2024) to non-linear settings, thus broadening its applicability.
@@
    We provide a curated set of benchmark problems specifically designed for Meta-Learning of dynamical systems. This collection encompasses a diverse range of problems frequently encountered in the physical science literature, all accessible through a unified interface: https://github.com/ddrous/gen-dynamics.
@@ 2 Related Work
While LEADS excels at interpolation tasks, it performs poorly during extrapolation Kirchmeyer et al. (2022). Furthermore, it requires retraining a new network $g^{e}_{\psi}$ each time a new environment is encountered, which can be constraining in scenarios where adaptation is frequently required. Before LEADS, other MTL approaches had been proposed outside the context of dynamical systems Caruana (1997); Rebuffi et al. (2017, 2018); Lee and Parish (2021). Still, they do not address the crucial adaptation to new tasks which is our focus.
@@
Arguably the most successful Meta-Learning method for physical systems is CoDA Kirchmeyer et al. (2022), which assumes that the underlying system is described by parametrized differential equations whose form is shared by all environments. However, these equations differ by the values of the vector field’s weights, which are produced by a (linear) hypernetwork. For the environment indexed by $e$, these weights are computed by [^1] The general CoDA formulation encompasses the GMBL adaptation rule Kirchmeyer et al. (2022). Furthermore, MTL models can be identified to Equation 3 with $\theta^{c}=\mathbf{0}$.
@@
where $\theta^{c}$ and $W$ are shared across environments, and $\xi^{e}\in\mathbb{R}^{d_{\xi}}$ is an environment-specific latent context vector (or simply context). While it achieves state-of-the-art performance on many physical systems, the main limitation of CoDA is its hypernetwork approach, which might hinder parallelism and scaling to large root or target networks. In practice, since they involve two networks instead of one, methods based on hypernetworks require more computational resources to train, and exhibit a more complex optimization landscape Chauhan et al. (2023).
@@
Recently, Blanke and Lelarge (2024) modeled the variability of linearly parameterized dynamical systems with an affine function of low-dimensional environment-specific context vectors. They empirically showed that this improved interpretability, generalization abilities, and computation speed. In our work, we similarly explain non-linear systems, thus generalizing existing work with higher-order Taylor expansion. Additionally, we extract additional benefits such as parallelizability and uncertainty quantification.
3 Neural Context Flow (NCF)
@@
We learn a single vector field for all environments in our training set $\operatorname{\mathcal{D}_{\text{tr}}}$. The same vector field will be reused, unchanged, for future testing and adaptation to environments in $\operatorname{\mathcal{D}_{\text{ad}}}$.
The vector field $f_{\theta}$ is assumed to not only be continuous, but also smooth in its second argument $\xi$. Exploiting this inductive bias, we “collect” information from other environments by Taylor-expanding $f_{\theta}$ around any other $\{\xi^{j}\}_{j\in\mathrm{P}}$, where $\mathrm{P}\subseteq\operatorname{[\![}1,m\operatorname{]\!]}$ is a context pool [^2] We note that $\mathrm{P}$ might include $e$ itself, and is reconstituted at each evaluation of loss function (9). Only its size $p$ is constant. containing $p\geq 1$ distinct environment indices. This gives rise, for fixed $e$ and $i$, to $p$ Neural ODEs
@@
where $o(\cdot)$ captures negligible residuals. Equation 6 directly consists of a Jacobian-Vector Product (JVP), making its implementation memory-efficient. Since higher-order Taylor expansions of vector-valued functions do not readily display the same property, we provide the following proposition to facilitate the second-order approximation.
Proposition 3.1 (Second-order Taylor expansion with JVPs).
@@
By setting $\xi^{e}:=\xi$, and $\xi^{j}:=\tilde{\xi}$, Proposition 3.1 yields an expression for $T^{2}_{f_{\theta}}(\cdot,\xi^{e},\xi^{j})$ in terms of JVPs, an implementation of which we detail in Appendix E. During training (as described below) trajectories from all $p$ Neural ODEs are used within the loss function.
Figure 2: Illustration of the Neural Context Flow (NCF). (a) Given an initial condition $x_{i}^{e}(0)$ for a training trajectory $i$ in the environment $e$, NCF predicts in parallel, several candidate trajectories $\{\hat{x}^{e,j}_{i}\}_{j\in\mathrm{P}}$ that are all compared to the ground truth $x_{i}^{e}$, upon which $\{\theta,\xi^{e}\}$ is updated. (b) Illustration of the 3-networks architecture for $f_{\theta}$ where the state vector and the arbitrary context are projected into the same representational space before they can interact inside a main network (in purple).
@@
We depict the NCF framework in Figure 2, along with a 3-networks architecture that projects the contexts and the state vectors into the same representational space before they can interact. Concretely, $x_{i}^{e}(t)$ and the context $\xi$ are first processed independently into $\tilde{x}_{i}^{e}(t)$ (by a state network in blue) and $\tilde{\xi}$ (by a context network in green) respectively, before they are concatenated and fed to a main network (in purple) to produce $f_{\theta}(x_{i}^{e}(t),\xi)$. This allows the framework to automatically balance the potentially nonlinear influence of the context with that of the state vector at each evaluation of the vector field.
@@ 3.1 Meta-Training
Starting from the same initial state $\hat{x}^{e,j}_{i}(0):=x^{e}_{i}(0)$, the Neural ODEs (5) are integrated using a differentiable numerical solver Kidger (2022); Nzoyem et al. (2023); Poli et al. ; Chen (2018):
@@
The resulting candidate trajectories are evaluated at time steps $\left\{t_{n}\right\}_{n\in\operatorname{[\![}1,N\operatorname{]\!]}}$ such that $t_{1}=0\text{ and }t_{N}=T$. We feed this to a supervised (inner) loss function
@@
where $d,d_{\xi},$ and $d_{\theta}$ are the dimensions of the state space, context vectors, and flattened network weights, respectively. $\|\cdot\|_{1}$ and $\|\cdot\|_{2}$ denote the $L^{1}$ and $L^{2}$ norms, with $\lambda_{1}$ and $\lambda_{2}$ their penalty coefficients, respectively. The $L^{1}$ regularization term is particularly vital in promoting sparsity in a possibly high-dimensional context vector. To ensure its independence from environment count, context pool size, and trajectory count per environment, the overall MSE loss function $\mathcal{L}$ is expressed as in Equation 10; after which it is minimized wrt both weights $\theta$ and contexts $\xi^{1:m}:=\{\xi^{e}\}_{e\in\operatorname{[\![}1,m\operatorname{]\!]}}$ via gradient descent, alternating between updates:
@@
The most effective way to train NCFs is via proximal alternating minimization as described in Algorithm 1. Although more computationally demanding compared to ordinary alternating minimization (see Algorithm 3), it is adept at dealing with non-smooth loss terms like the $L^{1}$ norm in Equation 9 Parikh et al. (2014). Not to mention that ordinary alternating minimization (Algorithm 3) can easily lead to sub-optimal convergence Attouch et al. (2010); Li et al. (2019), while its proximal counterpart converges, with random initialization, almost surely to second-order stationary points provided mild assumptions are satisfied (see Theorem 3.2). We provide a short discussion on those assumptions in Appendix A.
The above comparison demands the definition of two variants of Neural Context Flows. NCF-$\mathit{t}_{2}$ uses second-order Taylor expansion in (5), and is trained via proximal alternating minimization (Algorithm 1). NCF-$\mathit{t}_{1}$ on the other hand, is implemented using a first-order Taylor expansion, and trained using ordinary alternating minimization (Algorithm 3). Although less robust than NCF-$\mathit{t}_{2}$, NCF-$\mathit{t}_{1}$ is faster and serves as a powerful baseline in our experiments.
Theorem 3.2 (Convergence to second-order stationary points).
@@
The proof of Theorem 3.2 is straightforward by adapting Assumptions 1 and 4, then Theorem 2 of Li et al. (2019). ∎

### v5

@@ Abstract
Neural Ordinary Differential Equations (NODEs) often struggle to adapt to new dynamic behaviors caused by parameter changes in the underlying physical system, even when these dynamics are similar to previously observed behaviors. This problem becomes more challenging when the changing parameters are unobserved, meaning their value or influence cannot be directly measured when collecting data. To address this issue, we introduce Neural Context Flow (NCF), a robust and interpretable Meta-Learning framework that includes uncertainty estimation. NCF uses Taylor expansion to enable contextual self-modulation, allowing context vectors to influence dynamics from other domains while also modulating themselves. After establishing theoretical guarantees, we empirically test NCF and compare it to related adaptation methods. Our results show that NCF achieves state-of-the-art Out-of-Distribution performance on 5 out of 6 linear and non-linear benchmark problems. Through extensive experiments, we explore the flexible model architecture of NCF and the encoded representations within the learned context vectors. Our findings highlight the potential implications of NCF for foundational models in the physical sciences, offering a promising approach to improving the adaptability and generalization of NODEs in various scientific applications.
@@ 1 Introduction
A prototypical autonomous dynamical system describes the continuous change, through time $t\in\mathbb{R}^{+}$, of a quantity $x(t)\in\mathbb{R}^{d}$. Its dynamics are influenced by its parameters $c\in\mathbb{R}^{d_{c}}$, according to the (ordinary) differential equation
@@
Consider the challenge of reconstructing the mechanical motion of an undamped pendulum given limited data from two distinct environments: Earth and Mars. Disregarding the physical variations between these environments (e.g., gravitational constants, ambient temperatures, etc.), one might employ a One-For-All (OFA) approach to learn a single environment-agnostic vector field from all available data Yin et al. (2021a). This model would struggle to adequately fit such heterogeneous data and would face difficulties generalizing to data from novel environments, e.g., the Moon. Fig. 1 illustrates this issue, using gravity as the underlying physical parameter. Alternatively, learning individual vector fields for each environment with a One-Per-Env (OPE) approach would miss inter-domain commonalities, proving both time-intensive and inhibiting rapid adaptation to new environments Yin et al. (2021a). Given these constraints, it becomes imperative to develop a methodology that can effectively learn what to learn from the aggregate data while simultaneously accounting for the unique properties of each environment.
@@
In Scientific Machine Learning Cuomo et al. (2022) (SciML), the problem of generalization has largely been tackled by injecting domain knowledge. It is commonly understood that adding a term in Eq. 1 that captures as much of the dynamics as possible leads to lower evaluation losses Yin et al. (2021b). For such terms to be added, however, it is essential to have knowledge of the physical parameters that change, which may then either be directly estimated, or predicted by a neural network within the vector field Rackauckas et al. (2020). We are naturally left to wonder how to efficiently learn a generalizable dynamical system when such physics are absent.
Under constantly changing experimental conditions, two major obstacles to learning the parameter dependence of a vector field can be isolated: • (P1) Limited data – SciML models can be data-intensive Hey et al. (2020); Yin et al. (2021b), and limited data in each environment may not be enough to learn a vector field suitable for all environments; • (P2) Unobserved parameters – during the data collection and modeling processes, one might be unfamiliar with the basic physics of the system. Solving these two problems would contribute to the efficient generalization of the learned models, particularly to related but Out-of-Distribution (OoD) datasets. Fast OoD adaptation would massively reduce cost and complement recent efforts towards foundational models in the physical sciences Subramanian et al. (2024); McCabe et al. (2023); Herde et al. (2024).
Neural Ordinary Differential Equations (Neural ODEs) Chen et al. (2018); Weinan (2017); Haber & Ruthotto (2017) have emerged as a powerful backbone for learning ordinary, stochastic, and controlled differential equations Kidger et al. (2020). Trained using Differentiable Programming techniques Nzoyem et al. (2023); Ma et al. (2021); Rackauckas et al. (2020), they have demonstrated broad utility with outstanding results in areas like chemical engineering Owoyele & Pal (2022), geosciences Shen et al. (2023), and climate modeling Kochkov et al. (2024). In the increasingly popular SciML subfield of solving parametric PDEs Li et al. (2020); Takamoto et al. (2023); Li et al. (2023); Subramanian et al. (2024); Koupaï et al. (2024), Neural ODEs occupy a place of choice due to their flexibility and performance Yin et al. (2021a); Lee & Parish (2021); Kirchmeyer et al. (2022). That said, existing methods seeking to generalize Neural ODEs to various parameter-defined environments fail to leverage information from environments other than the ones of interest. Not to mention the pervasive lack of interpretability and ways of accounting for the model’s uncertainty.
@@
This work presents Neural Context Flows (NCFs), a novel approach for multi-environment generalization of dynamical systems based on Neural ODEs. By leveraging the regularity of the vector field with respect to unobserved parameters, NCFs parameterize an environment-agnostic vector field and environment-specific latent context vectors to modulate the vector field. The vector field is Taylor-expanded about these context vectors, effectively allowing information to flow between environments. Our contribution is threefold:
@@
    We introduce a Meta-Learning methodology for enhancing the generalizability of dynamical systems. Our approach effectively addresses problems (P1) and (P2), challenging the prevailing notion that standard Deep Learning techniques are inherently inefficient for Out-of-Distribution (OoD) adaptation Mouli et al. (2024).
@@
    We present an interpretable framework for Multi-Task representation learning, incorporating a straightforward method for uncertainty estimation. This work extends the emerging trend of explainable linearly parameterized physical systems Blanke & Lelarge (2024) to non-linear settings, thus broadening its applicability. For affine systems, we provide a lean proof for the identifiability of their underlying parameters.
@@
    We provide a curated set of benchmark problems specifically designed for Meta-Learning of dynamical systems. This collection encompasses a diverse range of problems frequently encountered in the physical science literature, all accessible through a unified interface.
@@ 2 Related Work
While LEADS excels at interpolation tasks, it performs poorly during extrapolation Kirchmeyer et al. (2022). Furthermore, it requires retraining a new network $g^{e}_{\psi}$ each time a new environment is encountered, which can be constraining in scenarios where adaptation is frequently required. Before LEADS, other MTL approaches had been proposed outside the context of dynamical systems Caruana (1997); Rebuffi et al. (2017; 2018); Lee & Parish (2021). Still, they do not address the crucial adaptation to new tasks which is our focus.
@@
Arguably the most successful Meta-Learning method for physical systems is CoDA Kirchmeyer et al. (2022), which assumes that the underlying system is described by parametrized differential equations whose form is shared by all environments. However, these equations differ by the values of the vector field’s weights, which are produced by a (linear) hypernetwork. For the environment indexed by $e$, these weights are computed by [^1] The general CoDA formulation encompasses the GMBL adaptation rule Kirchmeyer et al. (2022). Furthermore, MTL models can be identified to Eq. 3 with $\theta^{c}=\mathbf{0}$.
@@
where $\theta^{c}$ and $W$ are shared across environments, and $\xi^{e}\in\mathbb{R}^{d_{\xi}}$ is an environment-specific latent context vector (or simply context). While it achieves state-of-the-art performance on many physical systems, the main limitation of CoDA is its hypernetwork approach, which might hinder parallelism and memory scaling to large root or target networks. In practice, methods based on hypernetworks require more computational resources to backpropagate and train, and exhibit a more complex optimization landscape Chauhan et al. (2023).
@@
This local approximation strategy finds numerous applications in SciML, notably helping establish single-trajectory geometrical requirements for linear and affine system identification Duan et al. (2020). Recently, Blanke & Lelarge (2024) modeled the variability of linearly parameterized dynamical systems with an affine function of low-dimensional environment-specific context vectors. They empirically showed that this improved interpretability, generalization abilities, and computation speed. In our work, we similarly explain non-linear systems, thus generalizing existing work with higher-order Taylor expansion. Additionally, we extract benefits such as massive parallelizability and uncertainty estimation.
3 Neural Context Flow
@@
We learn a single vector field for all environments in our training set $\operatorname{\mathcal{D}_{\text{tr}}}$. The same vector field will be reused, unchanged, for future testing and adaptation to environments in a similarly-defined $\operatorname{\mathcal{D}_{\text{ad}}}$.
The vector field $f_{\theta}$ is assumed to not only be continuous, but also smooth in its second argument $\xi$. Exploiting this constraint, we “collect” information from other environments by Taylor-expanding $f_{\theta}$ around any other $\{\xi^{j}\}_{j\in\mathrm{P}}$, where $\mathrm{P}\subseteq\operatorname{[\![}1,m\operatorname{]\!]}$ is a context pool [^2] We note that $\mathrm{P}$ might include $e$ itself, and is reconstituted at each evaluation of Eq. 9. Its size $p$ is constant, each element indexing a distinct environment for computational efficiency (see Section A.6). containing $p:=|\mathrm{P}|\geq 1$ environment indices. This gives rise, for fixed $e$ and $i$, to $p$ Neural ODEs
@@
where $o(\cdot)$ captures negligible residuals. Eq. 6 directly consists of a Jacobian-Vector Product (JVP), making its implementation memory-efficient. Since higher-order Taylor expansions of vector-valued functions do not readily display the same property, we provide the following proposition to facilitate the second-order approximation.
Proposition 1 (Second-order Taylor expansion with JVPs).
@@
By setting $\xi^{e}:=\xi$, and $\xi^{j}:=\tilde{\xi}$, Proposition 1 yields an expression for $T^{2}_{f_{\theta}}(\cdot,\xi^{e},\xi^{j})$ in terms of JVPs, an implementation of which we detail in Appendix E. During training (as described below) trajectories from all $p$ Neural ODEs are used within the loss function.
Figure 2: (a) Illustration of the Neural Context Flow (NCF). Given an initial condition $x_{i}^{e}(0)$ for a training trajectory $i$ in the environment $e$, NCF predicts in parallel, several candidate trajectories $\{\hat{x}^{e,j}_{i}\}_{j\in\mathrm{P}}$ that are all compared to the ground truth $x_{i}^{e}$, upon which $\{\theta,\xi^{e}\}$ is updated. (b) Depiction of the 3-networks architecture for $f_{\theta}$, where the state vector and an arbitrary context vector are projected into the same representational space before they can interact inside the main network.
@@
We depict the NCF framework in Fig. 2, along with a 3-networks architecture that lifts the contexts and the state vectors into the same representational space before they can interact. Concretely, $x_{i}^{e}(t)$ and the context $\xi$ are first processed independently into $\tilde{x}_{i}^{e}(t)$ (by a state network in blue) and $\tilde{\xi}$ (by a context network in green) respectively, before they are concatenated and fed to a main network (in purple) to produce $f_{\theta}(x_{i}^{e}(t),\xi)$. This explicitely allows the model to account for the potentially nonlinear relationship between the context and the state vector at each evaluation of the vector field.
@@ 3.1 Meta-Training
Starting from the same initial state $\hat{x}^{e,j}_{i}(0):=x^{e}_{i}(0)$, the Neural ODEs in Eq. 5 are integrated using a differentiable numerical solver Kidger (2022); Nzoyem et al. (2023); Poli et al. ; Chen (2018):
@@
The resulting candidate trajectories are evaluated at specific time steps $\left\{t_{n}\right\}_{n\in\operatorname{[\![}1,N\operatorname{]\!]}}$ such that $t_{1}=0\text{ and }t_{N}=T$. We feed these to a supervised (inner) loss function
@@
where $d,d_{\xi},$ and $d_{\theta}$ are the dimensions of the state space, context vectors, and flattened network weights, respectively. $\|\cdot\|_{1}$ and $\|\cdot\|_{2}$ denote the regularizing $L^{1}$ and $L^{2}$ norms, with $\lambda_{1}$ and $\lambda_{2}$ their penalty coefficients, respectively. To ensure its independence from environment count, context pool size, and trajectory count per environment, the overall MSE loss function $\mathcal{L}$ is expressed as in Eq. 10; after which it is minimized wrt both weights $\theta$ and contexts $\xi^{1:m}:=\{\xi^{e}\}_{e\in\operatorname{[\![}1,m\operatorname{]\!]}}$ via gradient descent, alternating between updates:
@@
The most effective way to train NCFs is via proximal alternating minimization as described in Algorithm 1. Although more computationally demanding compared to ordinary alternating minimization (see Algorithm 3), it is adept at dealing with non-smooth loss terms like the $L^{1}$ norm in Eq. 9 Parikh et al. (2014). Not to mention that ordinary alternating minimization (Algorithm 3) can easily lead to sub-optimal convergence Attouch et al. (2010); Li et al. (2019), while its proximal counterpart converges, with random initialization, almost surely to second-order stationary points provided mild assumptions are satisfied (see Theorem 1). We provide a short discussion on those assumptions in Section A.3.
The above comparison demands the definition of two variants of Neural Context Flows. NCF-$\mathit{t}_{2}$ uses second-order Taylor expansion in equation 5, and is trained via proximal alternating minimization (Algorithm 1). NCF-$\mathit{t}_{1}$ on the other hand, is implemented using a first-order Taylor expansion, and trained using ordinary alternating minimization (Algorithm 3). Although less expressive than NCF-$\mathit{t}_{2}$, NCF-$\mathit{t}_{1}$ is faster and serves as a powerful baseline in our experiments.
Theorem 1 (Convergence to second-order stationary points).
@@
The proof of Theorem 1 is straightforward by adapting Assumptions 1 and 4, then Theorem 2 of Li et al. (2019). ∎